OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 625 and a3 = 25?

OpenStudy (campbell_st):

find the common ratio.... a2/625 = 25/a2 then the common ratio is 1/5 use r = 1/5 a = 625 \[T _{n} = ar^{n-1}\]

OpenStudy (campbell_st):

oops a = 625

OpenStudy (anonymous):

Thank you so much! (:

OpenStudy (anonymous):

Is it 0.04?